# Elastoresistance as probe of strain-controlled entropy from Kondo scattering

### v2

@@ Abstract
$\frac{\Delta R}{R_{0}}=\frac{\Delta\rho}{\rho_{0}}+(1+2\nu)\,\varepsilon$ (3)
@@
$m_{A_{1g}}=\frac{1}{1-\nu_{[100]}}\left[\frac{d(\Delta R/R_{0})_{[100]}}{d\varepsilon_{[100]}}+\frac{d(\Delta R/R_{0})_{[010]}}{d\varepsilon_{[100]}}\right]$ (4)
$m_{B_{1g}}=\frac{1}{1+\nu_{[100]}}\left[\frac{d(\Delta R/R_{0})_{[100]}}{d\varepsilon_{[100]}}-\frac{d(\Delta R/R_{0})_{[010]}}{d\varepsilon_{[100]}}\right]$ (5)
@@
$m_{A_{1g}}=\frac{1}{1-\nu_{[110]}}\left[\frac{d(\Delta R/R_{0})_{[110]}}{d\varepsilon_{[110]}}+\frac{d(\Delta R/R_{0})_{[\bar{1}10]}}{d\varepsilon_{[110]}}\right]$ (6)
$m_{B_{2g}}=\frac{1}{1+\nu_{[110]}}\left[\frac{d(\Delta R/R_{0})_{[110]}}{d\varepsilon_{[110]}}-\frac{d(\Delta R/R_{0})_{[\bar{1}10]}}{d\varepsilon_{[110]}}\right].$ (7)
Equations 4 and 6 are the relations providing two independent determinations of $m_{A_{1g}}$, serving as an internal consistency check on the analysis.
@@ II Effect of out-of-plane ($c$-axis) compression
$m_{A_{1g}}=\frac{\dfrac{d(\Delta R/R_{0})_{[100]}}{d\varepsilon_{[100]}}+\dfrac{d(\Delta R/R_{0})_{[010]}}{d\varepsilon_{[100]}}}{1-\nu_{[100]}-2\nu^{\prime}_{[100]}}$ (8)
@@ V In-plane thermal expansion measurement on thin YbRh2Si2 platelets
$\left(\frac{\Delta L}{L_{0}}\right)_{\mathrm{sample}}=\frac{\Delta L^{\mathrm{meas}}_{\mathrm{sample}}(T)-\Delta L_{\mathrm{empty\ cell}}(T)}{L_{0}}+\left(\frac{\Delta L}{L}\right)^{\mathrm{Cu}}_{\mathrm{lit}}(T),$ (9)
@@
$\alpha_{i}\sim\int\frac{1}{T}\frac{\partial(m_{i}R(T)/R_{\rm 300K})}{\partial T}dT+const.$ (10)

### v3

@@ Abstract
$\frac{\Delta R}{R_{0}}=\frac{\Delta\rho}{\rho_{0}}+(1+2\nu)\,\varepsilon$ (S1)
@@
$m_{A_{1g}}=\frac{1}{1-\nu_{[100]}}\left[\frac{d(\Delta R/R_{0})_{[100]}}{d\varepsilon_{[100]}}+\frac{d(\Delta R/R_{0})_{[010]}}{d\varepsilon_{[100]}}\right]$ (S2)
$m_{B_{1g}}=\frac{1}{1+\nu_{[100]}}\left[\frac{d(\Delta R/R_{0})_{[100]}}{d\varepsilon_{[100]}}-\frac{d(\Delta R/R_{0})_{[010]}}{d\varepsilon_{[100]}}\right]$ (S3)
@@
$m_{A_{1g}}=\frac{1}{1-\nu_{[110]}}\left[\frac{d(\Delta R/R_{0})_{[110]}}{d\varepsilon_{[110]}}+\frac{d(\Delta R/R_{0})_{[\bar{1}10]}}{d\varepsilon_{[110]}}\right]$ (S4)
$m_{B_{2g}}=\frac{1}{1+\nu_{[110]}}\left[\frac{d(\Delta R/R_{0})_{[110]}}{d\varepsilon_{[110]}}-\frac{d(\Delta R/R_{0})_{[\bar{1}10]}}{d\varepsilon_{[110]}}\right].$ (S5)
Equations S2 and S4 are the relations providing two independent determinations of $m_{A_{1g}}$, serving as an internal consistency check on the analysis.
@@ II Effect of out-of-plane ($c$-axis) compression
$m_{A_{1g}}=\frac{\dfrac{d(\Delta R/R_{0})_{[100]}}{d\varepsilon_{[100]}}+\dfrac{d(\Delta R/R_{0})_{[010]}}{d\varepsilon_{[100]}}}{1-\nu_{[100]}-2\nu^{\prime}_{[100]}}$ (S6)
@@ V In-plane thermal expansion measurement on thin YbRh2Si2 platelets
$\left(\frac{\Delta L}{L_{0}}\right)_{\mathrm{sample}}=\frac{\Delta L^{\mathrm{meas}}_{\mathrm{sample}}(T)-\Delta L_{\mathrm{empty\ cell}}(T)}{L_{0}}+\left(\frac{\Delta L}{L}\right)^{\mathrm{Cu}}_{\mathrm{lit}}(T),$ (S7)
@@
$\alpha_{i}\sim\int\frac{1}{T}\frac{\partial(m_{i}R(T)/R_{\rm 300K})}{\partial T}dT+const.$ (S8)